# Non-Gaussian Simultaneous Autoregressive Models with Missing Data

### v1

@@ 6 Conclusion
$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $=E_{f_{\boldsymbol{\delta}^{0}}}\Bigg{[}\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda_{s_{m}}}}\Big{(}\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S40)
$\qquad+\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\theta})\Big{)}\Bigg{]},$
@@
where, the random vector $\boldsymbol{\delta}^{0}$ has density $f_{\boldsymbol{\delta}^{0}}$, which follows a standard normal, and does not depend on $\boldsymbol{\lambda_{s_{m}}}$, and $t^{0}$ is the one-to-one vector-valued transformation from $\boldsymbol{\delta}^{0}=({\boldsymbol{\eta}^{0}}^{\top},{\boldsymbol{\epsilon}^{0}}^{\top})^{\top}$ to the parameter vector, such that $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$.
@@
see Poyiadjis et al., (2011). Substituting this expression into Equation (S40), and writing $E_{f_{\boldsymbol{\delta}}}(\cdot)$ for expectation with respect to $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ and because $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}_{o},\textbf{m}\mid\textbf{y}_{u},\boldsymbol{\xi},\boldsymbol{\psi})p(\textbf{y}_{u}\mid\boldsymbol{\xi})p(\boldsymbol{\xi},\boldsymbol{\psi})$, we get
$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $={E}_{f_{\boldsymbol{\delta}}}\Bigg{[}\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}\Big{(}\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S42)
@@
$={E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})^{\top}}{d\boldsymbol{\lambda}_{s_{m}}}\left(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right)\right].$
@@
The term $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\boldsymbol{\lambda_{s_{m}}}}$ in Equation (S42) is the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$ with respect to the variational parameters $\boldsymbol{\lambda}_{s_{m}}=(\boldsymbol{\mu}_{s_{m}}^{\top},\text{vech}(\textbf{B}_{s_{m}})^{\top},\textbf{d}_{s_{m}}^{\top})^{\top}$. We can express that $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})=\boldsymbol{\mu}_{s_{m}}+(\boldsymbol{\eta}^{0}\otimes\textbf{I}_{s_{m}})\text{vech}(\textbf{B}_{s_{m}})+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}$, where $\textbf{I}_{s_{m}}$ is the identity matrix of size $s_{m}$, and it can be further shown that $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}q^{0}_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi})=-(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}-\boldsymbol{\mu}_{s_{m}})$,
@@
$\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\boldsymbol{\mu}_{s_{m}}}=\textbf{I}_{s_{m}}\\ ~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}})}{d\text{vech}(\textbf{B}_{s_{m}})}={\boldsymbol{\eta}^{0}}^{\top}\otimes\textbf{I}_{s_{m}}.$ (S43)
@@
Then the derivatives of $\mathcal{L}(\boldsymbol{\lambda})$ with respect to $\boldsymbol{\mu}_{s_{m}},\textrm{vech}(\textbf{B}_{s_{m}})$, and $\textbf{d}_{s_{m}}$ are:
@@
$\begin{split}\nabla_{\boldsymbol{\mu}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~{}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})\\ &+(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}(\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0})],\end{split}$ (S44)
@@
$\begin{split}\nabla_{\textrm{vech}(\textbf{B}_{s_{m}})}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~{}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\eta}^{0}}^{\top}\\ &+(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}(\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\eta}^{0}}^{\top}],\end{split}$ (S45)
@@
$\begin{split}\nabla_{\textbf{d}_{s_{m}}}\mathcal{L}(\boldsymbol{\lambda})&=E_{f_{\boldsymbol{\delta}}}[\text{diag}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\epsilon}^{0}}^{\top}\\ &+(\textbf{B}_{s_{m}}\textbf{B}_{s_{m}}^{\top}+\textbf{D}_{s_{m}}^{2})^{-1}(\textbf{B}_{s_{m}}\boldsymbol{\eta}^{0}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\epsilon}^{0}}^{\top})].\end{split}$ (S46)
@@
The expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\mu}_{s_{m}}+\textbf{B}_{s_{m}}\boldsymbol{\eta}+\textbf{d}_{s_{m}}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})=\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~{}h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ in Equations (S44)-(S46) can be found in Section S4.3. The expectations in these gradients are estimated using a single sample of $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, which is drawn from $f_{\boldsymbol{\delta}^{0}}$, and $p(\textbf{y}_{u}\mid\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{o},\textbf{m})=p(\textbf{y}_{u}\mid t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}_{s_{m}}),\textbf{y}_{o},\textbf{m})$; see  Wijayawardhana et al., (2025) for further details.
S4.2 HVB-AllB algorithm
@@
This section presents the HVB-AllB algorithm for estimating SEMs with missing values, which is particularly suited for cases where both $n$ and $n_{u}$ are large.

### v2

@@ 9 Conclusion
$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $=\mathbb{E}_{f_{\boldsymbol{\delta}^{0}}}\Bigg[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}\Big(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S42)
$\qquad+\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~p(\textbf{y}_{o},\textbf{m}\mid\boldsymbol{\xi},\boldsymbol{\psi})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\Big)\Bigg],$
@@
where, the random vector $\boldsymbol{\delta}^{0}$ has density $f_{\boldsymbol{\delta}^{0}}$, which follows a standard normal, and does not depend on $\boldsymbol{\lambda}$, and $t^{0}$ is the one-to-one vector-valued transformation from $\boldsymbol{\delta}^{0}=({\boldsymbol{\eta}^{0}}^{\top},{\boldsymbol{\epsilon}^{0}}^{\top})^{\top}$} to the parameter vector, such that $(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}=t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$}.
@@
see Poyiadjis et al., (2011). Substituting this expression into Equation (S42), and writing $\mathbb{E}_{f_{\boldsymbol{\delta}}}[\cdot]$ for expectation with respect to $f_{\boldsymbol{\delta}}(\boldsymbol{\delta})=f_{\boldsymbol{\delta}^{0}}(\boldsymbol{\delta}^{0})p(\textbf{y}_{u}\mid\textbf{y}_{o},\textbf{m},\boldsymbol{\xi},\boldsymbol{\psi})$ and because $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}_{o},\textbf{m}\mid\textbf{y}_{u},\boldsymbol{\xi},\boldsymbol{\psi})p(\textbf{y}_{u}\mid\boldsymbol{\xi})p(\boldsymbol{\xi},\boldsymbol{\psi})$, we get
$\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ $=\mathbb{E}_{f_{\boldsymbol{\delta}}}\Bigg[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}\Big(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log p(\boldsymbol{\xi},\boldsymbol{\psi})$ (S44)
@@
$=\mathbb{E}_{f_{\boldsymbol{\delta}}}\left[\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})^{\top}}{d\boldsymbol{\lambda}}\left(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})-\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})\right)\right].$
S5.2 Derivation of $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$, $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\log q_{\boldsymbol{\lambda}}^{0}(\boldsymbol{\xi},\boldsymbol{\psi})$, and $\nabla_{\boldsymbol{\lambda}}\mathcal{L}(\boldsymbol{\lambda})$ for HVB
@@
In the reparameterisation gradient of the ELBO given in Equation (S44), $\frac{dt^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{d\boldsymbol{\lambda}}$ denotes the derivative of the transformation $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$} with respect to the variational parameters $\boldsymbol{\lambda}=(\boldsymbol{\mu}^{\top},\text{vech}(\textbf{B})^{\top},\textbf{d}^{\top})^{\top}$}. We can express that $t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})=\boldsymbol{\mu}+(\boldsymbol{\eta}^{0}\otimes\textbf{I}_{s})\text{vech}(\textbf{B})+\textbf{d}\circ\boldsymbol{\epsilon}^{0}$}, where $\textbf{I}_{s}$ is the identity matrix of size $s$. It can be further shown that $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~q^{0}_{\boldsymbol{\lambda}}(\boldsymbol{\xi},\boldsymbol{\psi})=-(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}((\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}-\boldsymbol{\mu})$,
@@
$\frac{\partial t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{\partial\boldsymbol{\mu}}=\textbf{I}_{s}\\ ,~~~~~~~~~~\frac{\partial t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{\partial\text{vech}(\textbf{B})}={\boldsymbol{\eta}^{0}}^{\top}\otimes\textbf{I}_{s}\\ ,~~~~~\text{and}~~~~~\frac{\partial t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda})}{\partial\textbf{d}}=\boldsymbol{\epsilon}^{0}.$ (S45)
@@
Then the derivatives of $\mathcal{L}(\boldsymbol{\lambda})$ with respect to $\boldsymbol{\mu},\textrm{vech}(\textbf{B})$, and d are:
@@
$\begin{split}\nabla_{\boldsymbol{\mu}}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0})],\end{split}$ (S46)
@@
$\begin{split}\nabla_{\textrm{vech}(\textbf{B})}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\delta}}}[\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}~\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\eta}^{0}}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\eta}^{0}}^{\top}],\end{split}$ (S47)
@@
$\begin{split}\nabla_{\textbf{d}}\mathcal{L}(\boldsymbol{\lambda})&=\mathbb{E}_{f_{\boldsymbol{\delta}}}[\text{diag}(\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi^{\top})^{\top}}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u}){\boldsymbol{\epsilon}^{0}}^{\top}\\ &+(\textbf{B}\textbf{B}^{\top}+\textbf{D}^{2})^{-1}(\textbf{B}\boldsymbol{\eta}^{0}+\textbf{d}\circ\boldsymbol{\epsilon}^{0}){\boldsymbol{\epsilon}^{0}}^{\top})].\end{split}$ (S48)
@@
The expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\mu}+\textbf{B}\boldsymbol{\eta}+\textbf{d}\circ\boldsymbol{\epsilon}^{0},\textbf{y}_{u})=\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$} in Equations (S46)-(S48) can be found in Section S5.3. The expectations in these gradients are estimated using a single sample of $\boldsymbol{\delta}=({\boldsymbol{\delta}^{0}}^{\top},\textbf{y}_{u}^{\top})^{\top}$, which is drawn from $f_{\boldsymbol{\delta}^{0}}$, and $p(\textbf{y}_{u}\mid\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{o},\textbf{m})=p(\textbf{y}_{u}\mid t^{0}(\boldsymbol{\delta}^{0},\boldsymbol{\lambda}),\textbf{y}_{o},\textbf{m})$; see  Wijayawardhana et al., (2025) for further details.
S5.3 Expressions for $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$
This section presents the derivatives (gradients) $\nabla_{(\boldsymbol{\xi}^{\top},\boldsymbol{\psi}^{\top})^{\top}}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$, which are required to compute the gradients of the reparameterised ELBO, defined in Equation (20) of the main paper.
S5.3.1 Gradients for SEM-Gau
For the SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$}, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho)^{\top}.$}
The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~p(\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\psi})$}. Note that, for $\sigma^{2}_{\textbf{e}}$, and $\rho$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S49)
where $\textbf{z}_{i}=({\textbf{x}^{*}_{i}}^{\top},y_{i})$, $\textbf{x}^{*}_{i}$$^{\top}$ is the $i^{\textrm{th}}$ row of matrix $\textbf{X}^{*}$, and $y_{i}$ is the $i^{\textrm{th}}$ element of the vector y; see Section 3 of the main paper.
S5.3.2 Gradients for SEM-t
For the SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$}, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\nu,\rho)^{\top}$.
The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\nu$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}\\ &+\sum_{i=1}^{n}m_{i}({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}},\end{split}$ (S51)
S5.3.3 Gradients for YJ-SEM-Gau
For the YJ-SEM-Gau with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$}, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\gamma)^{\top}$.
The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~p(\textbf{y}\mid\boldsymbol{\phi})+\textrm{log}~p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~p(\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\psi})$}. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, and $\gamma$, we utilise the transformations described in in Section S2.2, and place the priors on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}}-\frac{\boldsymbol{\psi}^{\top}\boldsymbol{\psi}}{2\sigma^{2}_{\boldsymbol{\psi}}},\end{split}$ (S52)
S5.3.4 Gradients for YJ-SEM-t
For the YJ-SEM-t with missing data, $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})=p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})p(\boldsymbol{\tau}\mid\boldsymbol{\phi})p(\boldsymbol{\phi})p(\boldsymbol{\psi})$}, where $\boldsymbol{\phi}=(\boldsymbol{\beta}^{\top},\sigma^{2}_{\textbf{e}},\rho,\nu,\gamma)^{\top}$.
The logarithm of $h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})$ is $\textrm{log}~p(\textbf{y}\mid\boldsymbol{\tau},\boldsymbol{\phi})+\textrm{log}~p(\textbf{m}\mid\textbf{y},\boldsymbol{\psi})+\textrm{log}~p(\boldsymbol{\tau}\mid\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\phi})+\textrm{log}~p(\boldsymbol{\psi})$. Note that, for $\sigma^{2}_{\textbf{e}}$, $\rho$, $\nu$, and $\gamma$, we utilise the transformations described in Section S2.2, and place the priors on the transformed parameters as in Table S3. This leads to
$\begin{split}\text{log}~h(\boldsymbol{\xi},\boldsymbol{\psi},\textbf{y}_{u})&\propto-\frac{n}{2}\omega^{\prime}+\frac{1}{2}\textrm{log}|\textbf{M}|-\frac{e^{-\omega{{}^{\prime}}}}{2}\textbf{r}^{\top}\textbf{M}\textbf{r}+\sum_{i=1}^{n}\text{log}\left(\frac{dt_{\gamma}(y_{i})}{dy_{i}}\right)\\ &+\sum_{i=1}^{n}m_{i}({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})-\text{log}(1+e^{({\textbf{x}^{*}_{i}}^{\top}\boldsymbol{\psi}_{\textbf{x}}+{y}_{i}\boldsymbol{\psi}_{y})})\\ &+\sum_{i=1}^{n}\textrm{log}~p(\tau_{i}\mid\nu)+\textrm{log}\left|\frac{\partial\tau_{i}}{\partial\tau_{i}^{\prime}}\right|-\frac{\boldsymbol{\beta}^{\top}\boldsymbol{\beta}}{2\sigma^{2}_{\boldsymbol{\beta}}}-\frac{{\omega^{\prime}}^{2}}{2\sigma^{2}_{\omega^{\prime}}}-\frac{{\rho^{\prime}}^{2}}{2\sigma^{2}_{\rho^{\prime}}}-\frac{{\nu^{\prime}}^{2}}{2\sigma^{2}_{\nu^{\prime}}}-\frac{{\gamma^{\prime}}^{2}}{2\sigma^{2}_{\gamma^{\prime}}},\end{split}$ (S53)
S5.4 HVB-AllB algorithm
@@
This section presents the MCMC steps used in the HVB-AllB algorithm for estimating SEMs with missing values, which is particularly suited for cases where both $n$ and $n_{u}$ are large.